OpenStudy (anonymous):

Two of the solutions for the equation 2x+3y=12 are (9 )and ( 4)

OpenStudy (hlambach):

Are you looking to see if those answers are correct?I don't really understand the question.

OpenStudy (anonymous):

Sry I'm looking for the other two that make the equation true

OpenStudy (hlambach):

Got it! Will be answering shortly. :)

OpenStudy (anonymous):

Thank you ^-^

OpenStudy (hlambach):

This is aa point on a line, right? So (9, 4), right?

OpenStudy (anonymous):

Yea

OpenStudy (hlambach):

But I don't even think the point (9, 4) is correct. See: 2(9) + 3(4) = 30 Are you sure you didn't type it wrong? Isn't it supposed to be (0, 4)?

OpenStudy (hlambach):

It's an easy typo to make, considering the 0 is right next to the 9 :)

OpenStudy (anonymous):

Yea Sry typo

OpenStudy (hlambach):

No problem! Okay, so what I always do to make the equation easier to solve is to put it in slope-intercept form: y = mx + b

OpenStudy (hlambach):

So just solve 2x + 3y = 12 for y. After that, graph the line and find out the second point